OpenStudy (anonymous):

(Trig) The London eye is 135m at it's highest point. At it's lowest point it is about 3m off the ground. It tales about 30 minutes to revolve entirely. Across the river, Big Ben is about 96m tall. If you were to take a ride on the London eye, for how long were you higher than Big Ben

OpenStudy (agl202):

Remove the 3 meters above the ground & that also includes the 3 meters less from Big Ben Your new height for Big Ben is 93 meters and diameter for the London Eye is 132 meters. Then find the speed as (implying constant) 1/30 rev/min = pi/15 rad/min as1 rev/min = 2*pi rad/min. Subtract 93 meters (Big Ben) - 66 meters (radius of London Eye) = 27 meters as you shall make a right triangle. adjacent = 27 meters; hypotenuse = 66 meters so the angle is arc cos(9/22). 2 * cos^-1(9/22) radians * (15/pi minute/radians) So, ur answer would make 30*cos^-1(9/22)/pi minutes.

pooja195 (pooja195):

Don't forget to give credit to your sources. https://answers.yahoo.com/question/index?qid=20140605214626AAuaUPB

OpenStudy (anonymous):

haha, interesting development, but an answer is an answer no matter where it is from.

pooja195 (pooja195):

Plagiarizing is never ok. And its definitely not ok to change a few words and call it their work.

OpenStudy (anonymous):

very true, and I'm glad you brought it to attention in this instance, thank you.

OpenStudy (agl202):

That's right. Thx for for doin it for me.